Another I need help on, I am soooo dumb!

If f(x) = x2 +5, find f(a +h) -f(a)
Possible answers:
2ah + h2 + 10
2ah + h2 + 5
2ah + h2
h2
Pleeeze show me how to solve this so I can apply it to the other questions
Thanks to anyone

Answer by stanbon(48516)

If f(x) = x^2 +5, find f(a +h) -f(a)
---------
Do each term separately, then subtract:
f(a+h) = (a+h)^2+5 = a^2+2ah+h^2+5
f(a) = a^2+5
-----------
Now subtract to get:
=2ah+h^2
---------
Cheers,
Stan H.